# CaMMT: Benchmarking Culturally Aware Multimodal Machine Translation

### v1

@@ Abstract
Translating cultural content poses challenges for machine translation systems due to the differences in conceptualizations between cultures, where language alone may fail to convey sufficient context to capture region-specific meanings. In this work, we investigate whether images can act as cultural context in multimodal translation. We introduce CaMMT, a human-curated benchmark of over 5,800 triples of images along with parallel captions in English and regional languages. Using this dataset, we evaluate five Vision Language Models (VLMs) in text-only and text+image settings. Through automatic and human evaluations, we find that visual context generally improves translation quality, especially in handling Culturally-Specific Items (CSIs), disambiguation, and correct gender usage. By releasing CaMMThttps://huggingface.co/datasets/villacu/cammt, we aim to support broader efforts in building and evaluating multimodal translation systems that are better aligned with cultural nuance and regional variation.
@@
Emilio Villa-Cueva✝,★, Sholpan Bolatzhanova✝,★, Diana Turmakhan✝,★, Kareem Elzeky, Henok Biadglign Ademtew, Alham Fikri Aji, Israel Abebe Azime, Jinheon Baek, Frederico Belcavello, Fermin Cristobal, Jan Christian Blaise Cruz, Mary Dabre, Raj Dabre, Toqeer Ehsan, Naome A Etori, Fauzan Farooqui, Jiahui Geng, Guido Ivetta, Thanmay Jayakumar, Soyeong Jeong, Zheng Wei Lim, Aishik Mandal, Sofía Martinelli, Mihail Minkov Mihaylov, Daniil Orel, Aniket Pramanick, Sukannya Purkayastha, Israfel Salazar, Haiyue Song, Tiago Timponi Torrent, Debela Desalegn Yadeta, Injy Hamed, Atnafu Lambebo Tonja, Thamar Solorio Core Authors (MBZUAI)
@@ 1 Introduction
Translation brings cultures into contact. It usually involves deciding how much of the foreignness to keep in the resulting translation, and invariably involves blending cultures to some degree Aixela (1999). As pointed out by Hershcovich et al. (2022), part of the difficulty arises from the different conceptualizations that each culture holds. Translators must, therefore, choose suitable strategies for adapting vocabulary as well as deciding whether to conserve or substitute foreign elements. Conforming the source text to the target culture by substituting unknown elements with familiar ones can ease comprehension, yet it simultaneously erases traces of the original culture Venuti (2003). Conversely, ignoring an adequate vocabulary choice that accounts for regional variation in the target language risks misinterpretation, as lexical choice directly shapes how readers understand a text Szymańska (2017).
@@
  • RQ4: Which translation strategies do native speakers prefer in the case of Culturally-Specific Items (CSIs)?
@@
  • Culturally-Specific MMT Dataset: We present CaMMT, a human-curated corpus of $5,817$ image-captions triples, where the captions are collected for both English and regional languages. The dataset spans $19$ languages and $23$ regions. Additionally, a dedicated dataset of over 1,550 image-captions tuples is collected, covering different translation strategies for CSIs [^2] We will make the corpus publicly available..
  • Insights into visual grounding for culture-aware translation: We evaluate five VLMs on CaMMT to assess the impact of visual grounding on human preferences and performance in automatic metrics. Through these experiments, we find that visual context improves translation outputs. Native speakers tend to prefer multimodal translations because they better preserve CSIs, resolve lexical ambiguities, and reflect correct gender usage, highlighting aspects of quality that standard metrics often fail to fully reflect.
@@ 2 Related Work
Recent benchmarks such as CVQA Romero et al. (2024), CulturalVQA Nayak et al. (2024), ALM-bench Vayani et al. (2025), and FoodieQA Li et al. (2024) demonstrate growing progress in regional image understanding within VLMs. However, none of these works study how imagery can affect translation across cultures.
Together, these studies motivate our evaluation on the multimodal translation ability of VLMs, where we introduce the first human-curated benchmark for culturally grounded MMT.
@@ 3 CaMMT Dataset
To ensure the correctness of the generated caption pairs, we involved native speakers (annotators) for each of the languages that participated in the original data curation and are co-authors of this paper. The annotators were asked to complete three tasks: (1) evaluate and ensure the grammatical correctness and parallelism of the generated pairs in English and regional language by correcting captions when needed, (2) ensure Culturally-Specific Items in regional language captions are preserved and (3) categorize each of the pairs into three categories: (a) Not culturally relevant sentences, (b) do not contain any Culturally-Specific Items (Non-CSI) or (c) contain Culturally-Specific Items (CSI).
We followed the work of Aixela (1999) to provide the annotators with a definition of CSIs. To achieve a better coverage of translation strategies for CSIs (as previously discussed in Section 2), we asked them to further categorize sentence pairs with CSI into (i) CSI with possible translation - captions containing CSIs that have culturally equivalent terms that can convey an equivalent meaning when translated into English and (ii) CSI forced translation - captions containing CSIs that do not have any equivalent translation in English. For each sentence containing CSI items, we asked the annotators to provide both conserved (retaining CSIs) and substituted (using familiar equivalents) English translations, then select their preferred version as native speakers.
For example, in the possible translation category, the Mexican term tianguis can be translated as flea market, as in: “The name for this type of Mexican informal market is tianguis” (conserved) or “The name for this type of Mexican informal market is flea market” (substituted). In contrast, a forced translation case is: “The name of the Egyptian food in the glass plate in the picture is Hawawshi” (conserved) and “The name of the Egyptian food in the glass plate in the picture is minced meat sandwich” (substituted), where the original term lacks an exact English equivalent. For forced translations, they provide the closest possible English approximation. We provide the annotation guidelines in Appendix G.
@@
In total, CaMMT comprises 23 regions with 19 different languages, with a total of 5,817 triples with additional 1,550 with conserved and substituted CSIs for targeted analysis. We present representative samples in Figure 1, and report the number of triples per language included in the corpus in Appendix A.
4 Methodology
This section outlines the experimental methodology, including model selection and evaluation setup of both human and automatic assessments used to measure translation quality across text-only and multimodal conditions.
4.1 VLMs for Multimodal Machine Translation
@@
To initially assess the ability of VLMs in grounding translations using images, we conduct a control experiment on the CoMMuTe dataset, comparing VLMs against strong task-specific MT and MMT baselines. In the text-only setting, models are prompted to translate from English to the target language. In the text+image setting, they are additionally provided with an image and prompted to use it as context for the translation (see Appendix D for prompt details). Importantly, no further instructions are given regarding the nature of the disambiguation task. We evaluate five VLMs: Gemma 3 27B and 12B Team et al. (2025), Qwen 2.5-VL 32B and 8B Bai et al. (2025), and Gemini 2.0 Flash Team et al. (2024).
Results presented in Table 1 demonstrate consistent and significant improvements in the performance of VLMs in the text+image over text-only setting. Moreover, the BLEU scores achieved by VLMs match or surpass those of NLLB-600M and NLLB-3.3B, strong baselines, as well as dedicated MMT systems. These findings confirm that VLMs can indeed leverage visual context to guide translation decisions. Based on this validation, we continue with VLMs as our testbeds to probe how visual grounding influences translation choices in our culturally relevant dataset, particularly in the handling of CSIs and region-sensitive translation.
4.2 Experimental Setup
We evaluate the five VLMs discussed in the previous section, which cover both closed- and open-weights models at different parameter scales, in both text-only and text+image setups. In the text+image setting, we do not instruct models to use images as a cultural reference, only as additional context, allowing us to observe their default effect in translation. To evaluate the impact of visual input on translation quality, we conduct both human preference evaluation and automatic evaluation using standard machine translation metrics.
@@ 4.2 Experimental Setup
Human Preference Evaluation
For 21 of the CaMMT regions, native speakers are presented with anonymized translations from three models—Qwen2.5-VL 32B, Gemma 3 27B, and Gemini 2.0 Flash—generated under both text-only and text+image settings. For each instance, they select the preferred translation and specify the reason for their preference from a predefined set: “CSI is preserved,” “Correct gender,” “Disambiguates word,” or “Regionally appropriate phrasing”. We identified this set of reasons based on an analysis carried out in preliminary experiments on a subset of languages. Annotators are also allowed to specify other reasons if none of the previous reasons explain the preference. In Appendix H, we present the instructions provided for this evaluation.
Automatic Evaluation
@@ 5.1 Effect of Visual Grounding
We base our main analyses on chrF++ as it has shown higher correlation with human judgments over BLEU Popović (2017); Kocmi et al. (2021). Figure 4 reports chrF++ for 23 regions across 19 language pairs. In the $X\to En$ direction, most regions show improvements with image-grounded translations, with a few exceptions (e.g., Japan, Indonesia, and China). In the $En\to X$ direction, the benefit of multimodality is less consistent: while Gemini demonstrates clear gains, other models show mixed trends, with no systematic advantage or degradation from adding images. We present the results on BLEU and BertScore in Appendix C, which reflect a similar pattern. Additionally, in Appendix E we report average chrF++ scores per CVQA-category.
@@
In Table 2, we report the average chrF++ scores across languages (for BLEU and BERT scores, refer to Appendix C). Notably, the addition of image context consistently improves performance across most VLMs, with gains most pronounced in the $X\to En$ direction. Both evaluations support the conclusion that visual grounding improves translation quality for most languages, particularly when translating from regional languages to English. For the reverse direction, benefits are model-specific: native speakers still tend to prefer image-grounded translations from open-weight models, but this is not always reflected in automatic metrics.
@@ 5.2 Reasons Behind Preferences
Beyond human preferences, we further analyze VLMs’ ability to handle CSIs in translation. Specifically, we compute the average proportion of translations in which a CSI is preserved when the source sentence contains a substituted ($En_{Sub}\to X$) or conserved ($En_{Cons}\to X$) CSI. To do this, we use GPT-4o in a two-step process: (1) extract the CSI from the conserved version of each sample, and (2) check whether it appears in the model-generated translation. Details of this procedure are provided in Appendix I. We then compute the percentage of CSI preservation by dividing the number of retained instances by the total number of samples.
Results presented in Table 4 show that, in the substituted setting, the inclusion of images leads to a higher rate of CSI preservation, indicating the model’s ability to retrieve appropriate region-specific concepts with visual grounding. In contrast, for the conserved setting, the image’s influence is less consistent, though it still appears to affect retention patterns to some extent.
@@
Table 4: Average percentage of preserved CSIs across languages. A value of 100 indicates that all CSIs are retained in the translation; 0 indicates none are preserved. Appendix I reports per-language differences and the average impact of images.
@@ 6 Discussion
Visual grounding generally improves translation quality, particularly in ways that are meaningful to human evaluators. While gains in automatic metrics such as BLEU and chrF++ may appear modest, human preference evaluations tell a richer story: In 85 out of 126 model–language–direction comparisons (67.5%), native speakers preferred multimodal translations, underscoring the value of images for improving cultural and semantic alignment of translations.
@@
That said, in 37 out of 126 comparisons (29.4%), text-only translations were preferred, indicating that visual input can occasionally degrade translation quality. Understanding why this occurs remains an open question and is an important direction for future work. Moreover, the relatively small gains in automatic metrics are consistent with patterns observed in earlier multimodal MT studies Futeral et al. (2022), underscoring the need for improved evaluation methods that more accurately reflect the contribution of visual context, particularly in multicultural scenarios.
Forced-C Forced-S Possible-C Possible-S
Latin 94$\pm$6.7 6$\pm$6.7 63$\pm$29.4 37$\pm$29.4
Non-Latin 75$\pm$36.2 25$\pm$36.2 53$\pm$20.2 47$\pm$20.2
@@
Contrary to the predominant research direction in NLP on substitution strategies for unfamiliar CSIs, our findings suggest that native speakers often prefer conservation, especially when no culturally equivalent term exists in English. This trend holds across both Latin and non-Latin scripts, although the latter group shows greater variability, possibly due to transliteration. When equivalents are available, preferences are more balanced, but still do not lean completely toward substitution. These results point to the importance of incorporating script-aware translation strategies regarding CSIs in future research, highlighting the need for MT systems to better align with native speaker preferences by adapting conservation and substitution choices to regional and linguistic contexts.

### v2

@@ Abstract
Translating cultural content poses challenges for machine translation systems due to the differences in conceptualizations between cultures, where language alone may fail to convey sufficient context to capture region-specific meanings. In this work, we investigate whether images can act as cultural context in multimodal translation. We introduce CaMMT, a human-curated benchmark of over 5,800 triples of images along with parallel captions in English and regional languages. Using this dataset, we evaluate five Vision Language Models (VLMs) in text-only and text+image settings. Through automatic and human evaluations, we find that visual context generally improves translation quality, especially in handling Culturally-Specific Items (CSIs), disambiguation, and correct gender marking. By releasing CaMMThttps://huggingface.co/datasets/villacu/cammt, our objective is to support broader efforts to build and evaluate multimodal translation systems that are better aligned with cultural nuance and regional variations.
@@
Emilio Villa-Cueva✝,★, Sholpan Bolatzhanova✝,★, Diana Turmakhan✝,★, Kareem Elzeky, Henok Biadglign Ademtew, Alham Fikri Aji, Vladimir Araujo, Israel Abebe Azime, Jinheon Baek, Frederico Belcavello, Fermin Cristobal, Jan Christian Blaise Cruz, Mary Dabre, Raj Dabre, Toqeer Ehsan, Naome A Etori, Fauzan Farooqui, Jiahui Geng, Guido Ivetta, Thanmay Jayakumar, Soyeong Jeong, Zheng Wei Lim, Aishik Mandal, Sofía Martinelli, Mihail Minkov Mihaylov, Daniil Orel, Aniket Pramanick, Sukannya Purkayastha, Israfel Salazar, Haiyue Song, Tiago Timponi Torrent, Debela Desalegn Yadeta, Injy Hamed, Atnafu Lambebo Tonja, Thamar Solorio Core Authors (MBZUAI)
@@ 1 Introduction
Translation brings cultures into contact. It usually involves deciding how much foreignness to keep in the resulting translation and invariably involves blending cultures to some extent Aixela (1999). As pointed out by Hershcovich et al. (2022), part of the difficulty in deciding the right level of culture blending during translation arises from the different conceptualizations that each culture holds. Translators must, therefore, choose suitable strategies for adapting vocabulary as well as deciding whether to conserve or substitute foreign elements. Conforming the source text to the target culture by substituting unknown elements with familiar ones can ease comprehension, yet it simultaneously erases traces of the original culture Venuti (2003). Conversely, ignoring an adequate vocabulary choice that accounts for regional variation in the target language risks misinterpretation, as lexical choice directly shapes how readers understand a text Szymańska (2017).
@@
  • Culturally-Specific MMT Dataset: We present CaMMT, a human-curated corpus of $5,817$ image-captions triples, where the captions are collected for both English and regional languages. For triples containing CSIs, we also provide a separate split with 1,550 samples, where each includes two English translations: one conserving the term and another substituting it.
  • Insights into visual grounding for culture-aware translation: We evaluate five VLMs on CaMMT to assess the impact of visual grounding on human preferences and performance in automatic metrics. Through these experiments, we find that visual context improves translation outputs. Native speakers tend to prefer multimodal translations because they better preserve CSIs, resolve lexical ambiguities, and reflect proper gender marking, highlighting aspects of translation quality ignored by standard evaluation metrics.
@@ 2 Related Work
Recent benchmarks such as CVQA Romero et al. (2024), CulturalVQA Nayak et al. (2024), ALM‑bench Vayani et al. (2025), and FoodieQA Li et al. (2024) demonstrate growing progress in regional image understanding within VLMs. However, none of these works study how imagery can affect translation across cultures. Together, these studies motivate our evaluation on the multimodal translation ability of VLMs.
@@ 3 CaMMT Dataset
No images were used in this process to ensure that the phrasing of these seed captions (later refined by annotators) was not influenced by them. The simplicity of the statements, combined with human curation, further reduced the risk of any bias from the language model.
@@
To ensure the correctness of the generated caption pairs, we involved native speakers (annotators) for each of the languages that participated in the original data curation and are co-authors of this paper. The annotators were asked to complete three tasks: (1) evaluate and ensure the grammatical correctness and parallelism of the generated pairs in English and regional language by correcting captions when needed, (2) ensure CSIs in regional language captions are preserved and (3) categorize each of the pairs into three categories: (a) Not culturally relevant sentences, (b) culturally relevant, but do not contain any CSI (Non-CSI) or (c) contain CSI.
We borrowed the definition of CSIs provided to annotators from Aixela (1999). To achieve a better coverage of translation strategies for CSIs (as previously discussed in Section 2), we asked them to further categorize sentence pairs marked as containing CSIs into (i) CSI with possible translation - captions containing CSIs that have culturally equivalent terms that can convey an equivalent meaning when translated into English and (ii) CSI forced translation - captions containing CSIs that do not have any equivalent translation in English. For each sentence containing CSIs, we asked the annotators to provide both conserved (retaining CSIs) and substituted (using familiar equivalents) English translations, then select their preferred version as native speakers.
For example, in the possible translation category, the Mexican term tianguis can be translated as flea market, as in: “The name for this type of Mexican informal market is tianguis” (conserved) or “The name for this type of Mexican informal market is flea market” (substituted). In contrast, a forced translation case is: “The name of the Egyptian food in the glass plate in the picture is Hawawshi” (conserved) and “The name of the Egyptian food in the glass plate in the picture is minced meat sandwich” (substituted), where the original term lacks an exact English equivalent. For forced translations, they provide the closest possible English approximation. We provide the annotation guidelines in Appendix A.8.
@@
In total, CaMMT comprises 23 regions with 19 different languages, with a total of 5,817 triples with additional 1,550 with conserved and substituted CSIs for targeted analysis. We present representative samples in Figure 1, and report the number of triples per language included in the corpus in Appendix A.1.
4 VLMs for Multimodal Machine Translation
This section explains our motivation for using VLMs as off-the-shelf MMT systems and our evaluation framework. We first present the selected models and validate their effectiveness for the task, followed by a description of our evaluation setup, which measures translation quality through human and automatic assessments in both text-only and text+image conditions in CaMMT.
@@ 4 VLMs for Multimodal Machine Translation
To initially assess the ability of VLMs in grounding translations using images, we conduct a control experiment on the CoMMuTe dataset Futeral et al. (2022), comparing them against strong task-specific MT and MMT baselines. CoMMuTe consists of English sentences with ambiguous terms paired with two images that lead to different translations (e.g., mole may refer to an animal or a skin mark). Thus, improvements when images are provided indicate that the model is effectively leveraging visual context to disambiguate the source sentence.
In the text-only setting, models are prompted to translate from English to the target language. In the text+image setting, they are additionally provided with an image and prompted to use it as context for the translation (see Appendix A.5 for prompt details). Importantly, no further instructions are given regarding the nature of the disambiguation task. We evaluate five VLMs: Gemma 3 27B and 12B Team et al. (2025), Qwen 2.5-VL 32B and 8B Bai et al. (2025), and Gemini 2.0 Flash Team et al. (2024).
Results presented in Table 1 demonstrate consistent and significant improvements in the performance of VLMs in the text+image over text-only setting. Moreover, the BLEU scores achieved by VLMs match or surpass those of NLLB-600M and NLLB-3.3B, strong baselines, as well as dedicated MMT systems. These results confirm that VLMs can indeed leverage visual context to guide translation decisions. Based on this validation, we continue with VLMs as our testbeds to probe how visual grounding influences translation choices in our culturally relevant dataset.
We focus on the Gemini, Gemma, and Qwen families, covering closed-weight and open-weight models at different scales. For completeness, we also report automatic evaluation results for Aya-Vision Cohere (2025) in Appendix A.3, consistent with the main models’ findings.
All models are evaluated in both text-only and text+image setups. In the text+image setting, we do not explicitly instruct models to use images as a cultural reference, only as additional context, allowing us to observe their default effect in translation. To evaluate the impact of visual input on translation quality, we conduct both human preference evaluation and automatic evaluation using standard machine translation metrics and the curated pairs as ground truth.
@@
Human Preference Evaluation Setup
For 21 of the CaMMT regions, native speakers are presented with anonymized translations from three models—Qwen2.5-VL 32B, Gemma 3 27B, and Gemini 2.0 Flash—generated under both text-only and text+image settings. For each instance, they select the preferred translation and specify the reason for their preference from a predefined set: “CSI is preserved,” “Correct gender,” “Disambiguates word,” or “Regionally appropriate phrasing”. We identified this set of reasons based on an analysis carried out in preliminary experiments on a subset of languages. Annotators are also allowed to specify other reasons if none of the previous reasons explain the preference. In Appendix A.9, we present the instructions provided for this evaluation.
Automatic Evaluation Setup
@@ 5.1 Effect of Visual Grounding
We base our main analyses on chrF++ as it has shown higher correlation with human judgments over BLEU Popović (2017); Kocmi et al. (2021). Figure 4 reports chrF++ for 23 regions across 19 language pairs. In the $X\to En$ direction, most regions show improvements with image-grounded translations, with a few exceptions (e.g., Japan, Indonesia, and China). In the $En\to X$ direction, the benefit of multimodality is less consistent: while Gemini demonstrates clear gains, other models show mixed trends, with no systematic advantage or degradation from adding images. We present the results on BLEU and BertScore in Appendix A.4, which reflect a similar pattern. Additionally, in Appendix A.6 we report average chrF++ scores per CVQA-category.
@@
In Table 2, we report the average chrF++ scores across languages (for BLEU and BERT scores, refer to Appendix A.4). Notably, the addition of image context consistently improves performance across most VLMs, with gains most pronounced in the $X\to En$ direction. Both evaluations support the conclusion that visual grounding improves translation quality for most languages, particularly when translating from regional languages to English. For the reverse direction, benefits are model-specific: native speakers still tend to prefer image-grounded translations from open-weight models, but this is not always reflected in automatic metrics.
To ensure that the observed gains are due to the images’ content and not the fact that we are simply providing an image, we carry out a control experiment. We replicate the evaluation of this section using randomly sampled images (i.e., images from other items in CaMMT) in the T+I setting. We find that this consistently hurts performance, with chrF++ scores dropping across most languages and models (see Figure 5). Therefore, these results confirm that the observed gains are due to the actual image content rather than simply providing any image. At the same time, they show that unrelated images can be harmful to multimodal translation.
Figure 5: Average chrF++ scores for text+image (T+I) and text-only (T) settings when the image in the T+I setting is randomly sampled. Unlike Figure 4, where the image corresponds to the translated sentence, here we observe that unrelated images consistently lower chrF++ scores compared to the text-only setting.
@@ 5.2 Reasons Behind Preferences
Beyond human preferences, we further analyze VLMs’ ability to handle CSIs in translation. Specifically, we compute the average proportion of translations in which a CSI is preserved when the English source sentence contains a substituted ($En_{Sub}\to X$) or conserved ($En_{Cons}\to X$) CSI.
In the substituted setting, the source includes a substituted term instead of the CSI. We compute the percentage of times the model “recovers” the original CSI with the help of the image instead of keeping the substituted term or replacing it by other equivalent. In the conserved setting, the source already contains the CSI, and we evaluate the percentage of times the model preserves it after translation. This analysis is independent of annotator preferences or ground-truth references, and instead probes how images bias models’ translation choices. As in prior experiments, models are not given explicit instructions about handling CSIs.
To do this, we use GPT-4o in a two-step process: (1) extract the CSI from the conserved version of each sample, and (2) check whether it appears in the model-generated translation. Details of this procedure are provided in Appendix A.10. We then compute the percentage of CSI preservation by dividing the number of retained instances by the total number of samples.
Results presented in Table 4 show that, in the substituted setting, the inclusion of images leads to a higher rate of CSI preservation, indicating the model’s ability to retrieve appropriate region-specific concepts with visual grounding. In the conserved setting, the effect of images is less consistent: while CSIs are often preserved, we observe that image grounding can also lead to modifications of the terms, resulting in a small decrease in the proportion of retained CSIs compared to the text-only setting. These results suggest that images can bias models to recover CSIs that were previously replaced in English, but may also introduce variability in CSIs when the CSI is already conserved in the source.
@@
Table 4: Average percentage of preserved CSIs across languages. A value of 100 indicates that all CSIs are retained in the translation; 0 indicates none are preserved. Appendix A.10 reports per-language differences and the average impact of images.
Forced-C Forced-S Possible-C Possible-S
[…]
@@ 6 Discussion
Visual grounding generally improves translation quality, particularly in ways that are meaningful to human evaluators. While gains in automatic metrics such as BLEU and chrF++ may appear modest, human preference evaluations tell a richer story: In 85 out of 126 model–language–direction comparisons (67.5%) where a preference was stated, native speakers preferred multimodal translations, underscoring the value of images for improving cultural and semantic alignment of translations.
@@
That said, in 37 out of 126 comparisons (29.4%), text-only translations were preferred, indicating that visual input can occasionally degrade translation quality. Understanding why this occurs remains an open question and is an important direction for future work. Moreover, the relatively small gains in automatic metrics are consistent with patterns observed in earlier multimodal MT studies Futeral et al. (2022), underscoring the need for improved evaluation methods that more accurately reflect the contribution of visual context, particularly in multicultural scenarios. Finally, we observed that unrelated images can negatively affect translation; therefore, future work should also study how to develop new models that can, on the fly, decide when visual context should influence translation to improve the robustness of these systems to noisy visual information.
@@
Contrary to the predominant research direction in NLP on substitution strategies for unfamiliar CSIs, our findings suggest that native speakers often prefer conservation, especially when no culturally equivalent term exists in English. This trend holds across both Latin and non-Latin scripts, although the latter group shows greater variability. When equivalents are available, preferences are more balanced, but still do not lean completely toward substitution. These results point to the importance of incorporating script-aware translation strategies regarding CSIs in future research, highlighting the need for MT systems to better align with native speaker preferences by adapting conservation and substitution choices to regional and linguistic contexts.